OpenStudy (studyx3):

Will medal and fan!! Factor (please explain as much as you can) 4x^2-22x+12=0

OpenStudy (studyx3):

How would you go about dividing 22 by 4?

OpenStudy (bonnieisflash1.0):

need help

OpenStudy (stefrheart):

Actually here is another way so that it is easier. What are numbers that could multiply to 6?

OpenStudy (bonnieisflash1.0):

all numbers

OpenStudy (studyx3):

3 and 2, 6 and 1

OpenStudy (bonnieisflash1.0):

x = -1/2 = -0.500 x = 6

OpenStudy (stefrheart):

Okay good and what are numbers that when they are multiplied, the sum is 12?

OpenStudy (studyx3):

Okay this is where I always get frustrated

OpenStudy (studyx3):

6 and 2, 3 and 4

OpenStudy (bonnieisflash1.0):

right

OpenStudy (bonnieisflash1.0):

here's two answers x = -1/2 = -0.500 x = 6

OpenStudy (bonnieisflash1.0):

my answer

OpenStudy (stefrheart):

sorry i meant 4 earlier, typo :3 but anywho so now you know the factors: Factors for 4: 1 & 4, 2 &2 Factors for 12: 1 & 12, 2&6, 3&4 Now we have to try it in different combinations to see if they will work out.

OpenStudy (skullpatrol):

$$\Huge 4x^2-22x+12$$ $$\Huge 2(2x^2-11x+6)$$ $$\Huge 2(2x-?)(x-?)$$

OpenStudy (studyx3):

So they have to add 4 and multiply to 12? @stefrheart

OpenStudy (skullpatrol):

Do you understand what I did above?

OpenStudy (stefrheart):

Look at what skullpatrol did above @studyx3

OpenStudy (studyx3):

yes @skullpatrol

OpenStudy (skullpatrol):

Good, now use FOIL with all the possible factors of 6 in the places of the question marks.

OpenStudy (studyx3):

okay where did the 6 come from? I'm sorry I'm not very good at this

OpenStudy (studyx3):

I FOILed it and got 4x^2-18x+36

OpenStudy (skullpatrol):

These are all the possible combinations you need to FOIL to see if you can find the one that will give you the correct linear term: $$\Huge -11x$$ $$\Huge (2x-1)(x-6)$$ $$\Huge (2x-6)(x-1)$$ $$\Huge (2x-2)(x-3)$$ $$\Huge (2x-3)(x-2)$$

OpenStudy (studyx3):

okay I'm beginning to understand now. Thank you so much for your help!!!

OpenStudy (skullpatrol):

Thanks for trying to understand :-)

OpenStudy (skullpatrol):

Please check to see if you have copied the question correctly.